# Scalable piecewise smoothing with BART

### v3

@@ C.2 Proof of Theorem 1
Up to this point, we have focused on single tree learners. We now shift our attention to forests for the next sequence of results. We define a tree ensemble =$\mathcal{E}=\left\{\mathcal{T}^{1},...,\mathcal{T}^{T}\right\}$ and denote the approximating ensemble by $\hat{\mathcal{E}}$, obtained by accumulating all the anisotropic k-d trees. Henceforth, we define $\mathcal{F}_{\mathcal{E}}$ as the set of functions $f_{\mathcal{E}}$ which satisfy:
$f_{\mathcal{E}}(\bm{x})=\sum_{t=1}^{T}\sum_{k=1}^{K_{t}}\beta_{tk}\cdot\varphi\left(\omega_{tk}^{\top}\bm{x}+b_{tk}\right)$
@@
Let $\hat{\mathcal{T}}$ be the $\mathcal{Z}$-tree partition defined in C4. Under assumptions (A5) and (A6),
@@
Next, we verify the three conditions in Ghosal and van der Vaart, (2007) involving prior concentration and metric entropy with respect to the empirical $L_{2}$-norm norm .$\|.\|_{n}$ defined as $\|f\|_{n}^{2}=n^{-1}\sum_{i=1}^{n}|f(x_{i})|^{2}$. These conditions are affirmatively tested in the following lemmas.
@@
$-\log\Pi\left(f\in\mathcal{F}_{\hat{\mathcal{E}}}:||f-f_{0,\hat{\mathcal{E}}}||_{n}\leq C\epsilon_{n}\right)\lesssim n\epsilon_{n}^{2}.$
@@
Assuming that $|\varphi|\leq C_{\varphi}$, observe that for two different ensembles ^$\hat{\mathcal{E}}_{1}$ and ^$\hat{\mathcal{E}}_{2}$:
$\left\|f_{\hat{\mathcal{E}}_{1}}-f_{\hat{\mathcal{E}}_{2}}\right\|_{\infty}$ $=\left\|\sum_{t=1}^{T}\sum_{k=1}^{K_{t}}\beta_{tk}^{(1)}\cdot\varphi(\omega_{tk}^{(1)\top}\bm{x}+b_{tk}^{(1)})-\sum_{t=1}^{T}\sum_{k=1}^{K_{t}}\beta_{tk}^{(2)}\cdot\varphi(\omega_{tk}^{(2)\top}\bm{x}+b_{tk}^{(2)})\right\|_{\infty}$
$\leq\left\|\sum_{t}\sum_{k}\beta_{tk}^{(1)}\cdot(\varphi(\omega_{tk}^{(1)\top}\bm{x}+b_{tk}^{(1)})-\varphi(\omega_{tk}^{(2)\top}\bm{x}+b_{tk}^{(2)}))+(\beta_{kt}^{(1)}-\beta_{kt}^{(2)})\cdot\varphi(\omega_{kt}^{(2)\top}\bm{x}+b_{kt}^{(2)})\right\|_{\infty}$
$\leq 2C_{\varphi}\left\|\sum_{t}\sum_{k}\beta_{tk}^{(1)}\right\|_{\infty}+C_{\varphi}\left\|\sum_{t}\sum_{k}(\beta_{kt}^{(1)}-\beta_{kt}^{(2)})\right\|_{\infty}$
$\leq 2\sqrt{K^{*}}C_{\varphi}||B_{1}||_{2}+C_{\varphi}\sqrt{K^{*}}||B_{1}-B_{2}||_{2}$ (C5)
where $B_{l}$ is defined to be the flattened vector $\left(\beta_{11}^{(l)},...,\beta_{1,K_{1}}^{(l)},...,\beta_{T1}^{(l)},...,\beta_{TK_{T}}^{(l)}\right)$; where $l=1,2$, each of dimension $\sum_{t=1}^{T}K_{t}=K^{*}$. The last inequality follows by Cauchy-Schwarz on $L_{1}$ norm. Now, define a matrix $M^{K^{*}\times K^{*}}$ such that $MB$ is a product of independent standard normal priors. It is worthwhile to note that:
$\Pi(f\in\mathcal{F}_{\hat{\mathcal{E}}}:||f-f_{0,\hat{\mathcal{E}},\hat{B}}||_{\infty}\leq C\epsilon_{n})$ $\geq\Pi\left(B\in\mathbb{R}^{K^{*}}:||B-\hat{B}||_{2}\leq\frac{C_{1}\epsilon_{n}}{\sqrt{K^{*}}}-2||\hat{B}||_{2}\right)$
$=\Pi\left(B\in\mathbb{R}^{K^{*}}:||M(B-\hat{B})||_{2}\leq\frac{||M^{-1}||_{\text{spec}}C_{1}\epsilon_{n}}{\sqrt{K^{*}}}-2||M^{-1}||_{\text{spec}}||\hat{B}||_{2}\right)$
$\geq 2^{\frac{K^{*}}{2}}e^{-||M\hat{B}||_{2}^{2}}\Pi\left(B\in\mathbb{R}^{K^{*}}:||MB||_{2}\leq||M^{-1}||_{\text{spec}}\left(\frac{C_{1}\epsilon_{n}}{\sqrt{2K^{*}}}-\frac{||\hat{B}||_{2}}{\sqrt{2}}\right)\right)$
Here, spec denotes the spectral norm of a matrix. The last inequality follows from the arguments in Ghosal and van der Vaart, (2007, pg. 217). The induced prior for $||MB||_{2}^{2}\sim\chi^{2}_{K^{*}}$. As the spectral norm is assumed to be bounded, the quantity
$v_{n}=\left(\frac{\epsilon_{n}}{\sqrt{K^{*}}}-2||\hat{B}||_{2}\right)||M^{-1}||_{\text{spec}}<\frac{\epsilon_{n}}{\sqrt{K^{*}}}||M^{-1}||_{\text{spec}}\lesssim\epsilon_{n}$
Following the same arguments as Jeong and Ročková, (2023),
$\Pi\left(B\in\mathbb{R}^{K^{*}}:||MB||_{2}\leq\frac{C_{1}v_{n}}{\sqrt{2}}\right)\geq\frac{2}{K^{*}2^{K^{*}}\Gamma(\frac{K^{*}}{2})}(C_{1}v_{n})^{K^{*}}\exp\left(-C_{1}^{2}\frac{v_{n}^{2}}{4}\right)$
The logarithm of the RHS is bounded below by a constant multiple of $-(K^{*}+T)\log n-v_{n}^{2}\gtrsim-n\epsilon_{n}^{2}$. See that $\exp(-||M\hat{B}||_{2}^{2})$ is bounded owing to the following two inequalities:
@@
$||M\hat{B}||_{2}^{2}\leq||M||_{\text{spec}}^{2}||\hat{B}||_{2}^{2}\leq||M||_{\text{spec}}^{2}K^{*}||\hat{B}||_{\infty}^{2}\lesssim K^{*}\log n$
under the assumption (A1). ∎
@@
Let $\bar{K}_{n}\asymp\frac{n\epsilon_{n}^{2}}{\log n}$ and define the function spaces $\mathcal{F}_{\mathcal{E},M}^{(1)}=\{f\in\mathcal{F}_{\mathcal{E}}:||B||_{\infty}\leq M\}$ and
@@
$\log\mathcal{N}(\epsilon_{n},\mathcal{F}_{\bar{K}_{n},n^{C}}^{(1)},||.||_{n})\lesssim n\epsilon_{n}^{2}$
@@
Observe the following from the arguments provided by Jeong and Ročková, (2023), for $\mathscr{E}$ – the collection of all ensembles $\mathcal{E}$ containing partitions of size $K_{t}$:
$N(\epsilon_{n},\mathcal{F}_{\bar{K}_{n},n^{C}}^{(1)},||.||_{\infty})\leq\sum_{K^{t}\leq\bar{K}_{n}}\sum_{\mathcal{E}\in\mathscr{E}}N(\epsilon_{n},\mathcal{F}_{\mathcal{E},n^{C}}^{(1)},||.||_{\infty})$
The next steps can be seen as:
$\left\|f_{\hat{\mathcal{E}}_{1}}-f_{\hat{\mathcal{E}}_{2}}\right\|_{\infty}$ $=\sup_{\mathbf{x}\in[0,1]^{p}}=\left\|\sum_{t=1}^{T}\sum_{k=1}^{K_{t}}\beta_{tk}^{(1)}\cdot\varphi(\omega_{tk}^{(1)\top}\bm{x}+b_{tk}^{(1)})-\sum_{t=1}^{T}\sum_{k=1}^{K_{t}}\beta_{tk}^{(2)}\cdot\varphi(\omega_{tk}^{(2)\top}\bm{x}+b_{tk}^{(2)})\right\|_{\infty}$
[…]
Observe that $\mathscr{E}=\prod_{t=1}^{T}|\mathcal{P}_{K^{t},\mathcal{Z}}|\leq|\mathcal{P}_{\bar{K}_{n},\mathcal{Z}}|^{T}$, denoting the cardinality by $\mathcal{P}$. We can argue that
$\mathcal{N}\left(\epsilon_{n},\mathcal{F}_{\bar{K}_{n},n^{C}}^{(1)},||.||_{n}\right)\leq|\mathcal{P}_{\bar{K}_{n},\mathcal{Z}}|^{T}\bar{K}_{n}^{T}\mathcal{N}\left(\frac{\epsilon_{n}}{T\bar{K}_{n}C_{\omega}},\{B\in\mathbb{R}^{T\bar{K}_{n}}:||B||_{\infty}\leq n^{C}\},||.||_{\infty}\right)$
See that,
$\mathcal{N}\left(\frac{\epsilon_{n}}{T\bar{K}_{n}},\{\|B\|_{\infty}\leq n^{C}\},\|\cdot\|_{\infty}\right)\leq\left(\frac{6T\bar{K}_{n}K^{*}C_{\omega}n^{C}}{\epsilon_{n}}\right),$
This is obtained partially from Jeong and Ročková, (2023) that if $||f_{\mathcal{E}_{1}}-f_{\mathcal{E}_{2}}||_{\infty}\leq K^{*}||B_{1}-B_{2}||_{\infty}$, then we would have obtained
$\mathcal{N}\left(\frac{\epsilon_{n}}{T\bar{K}_{n}},\{\|B\|_{\infty}\leq n^{C}\},\|\cdot\|_{\infty}\right)\leq\left(\frac{3T\bar{K}_{n}n^{C}}{\epsilon_{n}}\right),$
However, since we have the upper bound involving both $||B_{1}-B_{2}||_{\infty}$ and $||B_{1}||_{\infty}$, the covering number must reflect that. To adjust the covering number, we need to potentially inflate the radius of the covering balls i.e., the radius $\nicefrac{{\epsilon_{n}}}{{T\bar{K}_{n}}}$ will effectively become smaller $\nicefrac{{\epsilon_{n}}}{{T\bar{K}_{n}(2K^{*}C_{\omega})}}$.
The proof concludes by observing the following inequality:
$|\mathcal{P}_{\bar{K}_{n},\mathcal{Z}}|^{T}\leq p^{T\bar{K}_{n}+p+1}(\max_{j}b_{j})^{T\bar{K}_{n}}$
The logarithm of this expression is $\lesssim n\epsilon_{n}^{2}$ under the assumption (A5). ∎
@@
$B_{n}=\left\{(f,\sigma):\sum_{i=1}^{n}K(p_{o,i},p_{f,\sigma,i})\leq n\epsilon_{n}^{2},\ \sum_{i=1}^{n}V(p_{0,i},p_{f,\sigma,i})\leq n\epsilon_{n}^{2}\right\}$

### v4

@@ C.2 Proof of Theorem 1
Up to this point, we have focused on single tree learners. We now shift our attention to forests for the next sequence of results. We define a tree ensemble $\mathcal{E}=\left\{\mathcal{T}^{1},...,\mathcal{T}^{T}\right\}$ and denote the approximating ensemble by $\hat{\mathcal{E}}$, obtained by accumulating all the anisotropic k-d trees. Henceforth, we define $\mathcal{F}_{\mathcal{E}}$ as the set of functions $f_{\mathcal{E}}$ which satisfy:
$f_{\mathcal{E}}(\bm{x})=\sum_{t=1}^{T}\sum_{k=1}^{K_{t}}\sum_{d=1}^{D}\beta_{tk,d}\cdot\varphi\left(\omega_{tk,d}^{\top}\bm{x}+b_{tk,d}\right)$
@@
Let $\hat{\mathcal{T}}$ be the $\mathcal{Z}$-tree partition defined in C3. Under assumptions (A5) and (A6),
@@
Next, we verify the three conditions in Ghosal and van der Vaart, (2007) involving prior concentration and metric entropy with respect to the empirical $L_{2}$-norm norm $\|.\|_{n}$ defined as $\|f\|_{n}^{2}=n^{-1}\sum_{i=1}^{n}|f(x_{i})|^{2}$. These conditions are affirmatively tested in the following lemmas.
@@
$\log\Pi\left(f\in\mathcal{F}_{\hat{\mathcal{E}}}:||f-f_{0,\hat{\mathcal{E}}}||_{n}\leq C\epsilon_{n}\right)\lesssim n\epsilon_{n}^{2}.$
@@
Let the fixed approximating ensemble be $\hat{\mathcal{E}}$, which consists of $T$ trees, where tree $t$ has $K_{t}$ leaves. Let $f_{0,\hat{\mathcal{E}}}$ be the best approximant from Lemma 1, and let $f$ be a function drawn from the prior over this ensemble. To clarify the notation for the outer-weight vectors: $B$ is the random vector of all outer weights $\{\beta_{tk,d}\}$ for the function $f$. It is constructed by flattening all weight vectors from all leaves of all trees into a single vector. Its total dimension is $K^{*}=D\times\sum_{t=1}^{T}K_{t}$, where $D$ is the number of ridge functions per leaf. This is the vector over which the Gaussian prior is placed, $B\sim\mathcal{N}_{K^{*}}(\bm{0},\Sigma_{\beta})$. Similarly, $\hat{B}$ is the fixed, non-random vector of outer weights corresponding to the specific best approximating function $f_{0,\hat{\mathcal{E}}}$. It has the same dimension and structure as $B$ and serves as the fixed “target” in the parameter space that our prior needs to concentrate around.
@@
Since Remark C1 has already established the fact that $|\varphi|<C_{\varphi},$ we can now proceed to bound the functional distance. Let $B_{1}$ and $B_{2}$ be two outer-weight vectors.
$\left\|f_{B_{1}}-f_{B_{2}}\right\|_{\infty}$ $=\left\|\sum_{t=1}^{T}\sum_{k=1}^{K^{t}}\sum_{d=1}^{D}(\beta_{tk,d}^{(1)}-\beta_{tk,d}^{(2)})\varphi(\hat{\bm{\omega}}_{tk,d}^{\top}\bm{x}+\hat{b}_{tk,d})\right\|_{\infty}$
$\leq\sum_{t=1}^{T}\sum_{k=1}^{K^{t}}\sum_{d=1}^{D}|\beta_{tk,d}^{(1)}-\beta_{tk,d}^{(2)}|\cdot|\varphi(\hat{\bm{\omega}}_{tk,d}^{\top}\bm{x}+\hat{b}_{tk,d})|$ (Triangle Ineq.)
$\leq C_{\varphi}\sum_{t=1}^{T}\sum_{k=1}^{K^{t}}\sum_{d=1}^{D}|\beta_{tk,d}^{(1)}-\beta_{tk,d}^{(2)}|=C_{\varphi}\|B_{1}-B_{2}\|_{1}$ $\text{(Since }|\varphi|\leq C_{\varphi}\text{)}$
$\leq C_{\varphi}\sqrt{K^{*}}\|B_{1}-B_{2}\|_{2}$ (Cauchy-Schwarz)
This links the functional distance with the Euclidean distance of the outer-weight vectors.
We want the event $\|f_{B}-f_{0,\hat{\mathcal{E}}}\|_{\infty}\leq C\epsilon_{n}$. This is guaranteed if $C_{\varphi}\sqrt{K^{*}}\|B-\hat{B}\|_{2}\leq C\epsilon_{n}$, using our derived bound with $B_{1}=B$ and $B_{2}=\hat{B}$. This defines a ball in the space of $B$ with radius $r_{n}\coloneqq\frac{C\epsilon_{n}}{C_{\varphi}\sqrt{K^{*}}}$. To find the prior mass, we map this to the standard normal space.
Let $Z=L^{-1}B\sim\mathcal{N}_{K^{*}}(\bm{0},I_{K^{*}})$ and $\zeta=L^{-1}\hat{B}$, where $\Sigma_{\beta}=LL^{\top}$. We know $\|B-\hat{B}\|_{2}=\|L(Z-\zeta)\|_{2}\leq\|L\|_{\text{spec}}\|Z-\zeta\|_{2}$. From (P2), $\|L\|_{\text{spec}}\leq C_{\beta}^{1/2}$. Thus, a sufficient condition to guarantee $\|B-\hat{B}\|_{2}\leq r_{n}$ is $C_{\beta}^{1/2}\|Z-\zeta\|_{2}\leq r_{n}$, which is $\|Z-\zeta\|_{2}\leq r_{n}/C_{\beta}^{1/2}$. As this defines a smaller volume, we have the lower bound:
$\Pi(\|B-\hat{B}\|_{2}\leq r_{n})\geq\Pi\left(\|Z-\zeta\|_{2}\leq\frac{r_{n}}{C_{\beta}^{1/2}}\right)$
We use a standard Gaussian concentration inequality (Ghosal and van der Vaart,, 2007, pg. 217): $P(\|Z-\zeta\|_{2}\leq r^{\prime})\geq e^{-\|\zeta\|_{2}^{2}/2}P(\|Z\|_{2}\leq r^{\prime})$. Here $r^{\prime}=r_{n}/C_{\beta}^{1/2}$. From (A8), we have $\|\zeta\|_{2}^{2}=\|L^{-1}\hat{B}\|_{2}^{2}\lesssim K^{*}$. For large $K^{*}$, bounds on the norm of a Gaussian vector give:
$P(\|Z\|_{2}\leq r^{\prime})\geq\left(\frac{(r^{\prime})^{2}}{K^{*}}\right)^{K^{*}/2}e^{-((r^{\prime})^{2}-K^{*})/2}$
Taking the negative log of the prior probability, the dominant term is $K^{*}\log(K^{*})$. Given that optimal balancing requires $K^{*}\asymp n\epsilon_{n}^{2}/\log n$, we have:
$K^{*}\log(K^{*})\lesssim\frac{n\epsilon_{n}^{2}}{\log n}\log\left(\frac{n\epsilon_{n}^{2}}{\log n}\right)\lesssim\frac{n\epsilon_{n}^{2}}{\log n}(\log n)=n\epsilon_{n}^{2}$
The final inequality follows by noting that $\log\left(\frac{n\epsilon_{n}^{2}}{\log n}\right)=\log(n\epsilon_{n}^{2})-\log(\log n)\asymp\frac{p}{2\bar{\alpha}+p}\log n-\log(\log n)\lesssim\log n$. Therefore, we conclude that $\log\Pi\left(\|f-f_{0,\hat{\mathcal{E}}}\|_{n}\leq C\epsilon_{n}\right)\lesssim n\epsilon_{n}^{2}$. ∎
@@
[…]
@@
$\log N(\epsilon_{n},\mathcal{F}_{\bar{K}_{n},n^{C}}^{(1)},||.||_{n})\lesssim n\epsilon_{n}^{2}$
@@
Because $\|g\|_{n}\leq\|g\|_{\infty}$ for every function $g$, any $\varepsilon_{n}$–cover in $\|\cdot\|_{\infty}$ is automatically an $\varepsilon_{n}$–cover in $\|\cdot\|_{n}$. Hence, we bound the covering number for the infinity norm, $\log N\bigl{(}\varepsilon_{n},\mathcal{F}_{\bar{K}_{n},\,n^{C}},\|\cdot\|_{\infty}\bigr{)}.$ The total covering number is bounded by the product of the number of possible model structures and the covering number for the parameters on a fixed structure.
First, we count the number of admissible ensembles, $|\mathscr{E}|$. An ensemble $\mathcal{E}$ consists of $T$ trees. A single tree with $K$ leaves has K1K-1 internal split nodes. Fix a split-net $\mathcal{Z}=\{\mathcal{Z}_{j}\}_{j=1}^{p}$ with $b_{\max}:=\max_{j}b_{j}(\mathcal{Z})$. The number of ways to form a single tree with $K$ leaves, $|\mathcal{P}_{K,\mathcal{Z}}|$, is bounded by:
$|\mathcal{P}_{K,\mathcal{Z}}|\leq\bigl{(}p\,b_{\max}\bigr{)}^{K-1}$
Using Assumption (A6), where $b_{\max}\lesssim n^{c_{b}}$, the log-cardinality is:
$\log|\mathcal{P}_{K,\mathcal{Z}}|\leq(K-1)(\log p+c_{b}\log n)\lesssim K\log n$
Our sieve $\mathcal{F}_{\bar{K}_{n},\,n^{C}}$ is a union over ensembles where each tree has at most $\bar{K}_{n}$ leaves. The total number of such ensemble structures is bounded by:
$\log|\mathscr{E}|\leq\sum_{t=1}^{T}\log|\mathcal{P}_{\bar{K}_{n},\mathcal{Z}}|\leq T\,\bar{K}_{n}\,\log n$
By the definition of the sieve, we choose $\bar{K}_{n}\asymp n\varepsilon_{n}^{2}/\log n$ to balance the model complexity with the approximation error. This yields:
$\log|\mathscr{E}|\lesssim T\left(\frac{n\varepsilon_{n}^{2}}{\log n}\right)\log n=T\cdot n\varepsilon_{n}^{2}\lesssim n\varepsilon_{n}^{2}$
assuming $T$ is a fixed constant.
Now, fix a single ensemble structure $\mathcal{E}$. This means the tree partitions and the inner weights $\{(\bm{\omega}_{tk,d},\bm{b}_{tk,d})\}$ are deterministic. We only need to cover the space of the outer-weight vectors $B$. The relationship between the functional distance and the parameter distance is:
$\bigl{\|}f_{\mathcal{E},B_{1}}-f_{\mathcal{E},B_{2}}\bigr{\|}_{\infty}$ $=\left\|\sum_{t=1}^{T}\sum_{k=1}^{K^{t}}\sum_{d=1}^{D}(\beta_{tk,d}^{(1)}-\beta_{tk,d}^{(2)})\varphi(\bm{\omega}_{tk,d}^{\top}\bm{x}+b_{tk,d})\right\|_{\infty}$
$\leq C_{\varphi}\|B_{1}-B_{2}\|_{1}\leq C_{\varphi}\,K^{*}\,\|B_{1}-B_{2}\|_{\infty}$
where $K^{*}=D\sum K^{t}$ is the total number of outer-weight parameters.
To achieve $\varepsilon_{n}$–accuracy, it suffices to cover the parameter cube $n^{C},n^{C}]^{K^{*}}$ with $\ell_{\infty}$–balls of radius $\delta_{n}=\frac{\varepsilon_{n}}{C_{\varphi}\,K^{*}}.$ The number of such balls, $N_{B}$, is given by:
$N_{B}=\left(\frac{2n^{C}}{\delta_{n}}\right)^{K^{*}}=\left(\frac{2C_{\varphi}K^{*}n^{C}}{\varepsilon_{n}}\right)^{K^{*}}$
Taking the logarithm gives $\log N_{B}\lesssim K^{*}\log n$. Since $K^{*}=D\sum_{t=1}^{T}K^{t}\leq D\cdot T\cdot\bar{K}_{n}$, and using the definition of $\bar{K}_{n}$:
$\log N_{B}\lesssim(D\cdot T\cdot\bar{K}_{n})\log n\asymp D\cdot T\cdot\left(\frac{n\varepsilon_{n}^{2}}{\log n}\right)\log n=D\cdot T\cdot n\varepsilon_{n}^{2}\lesssim n\varepsilon_{n}^{2}$
The total log-covering number is bounded by the sum of the log-cardinality of structures and the log-covering number of the parameters:
$\log N\bigl{(}\varepsilon_{n},\mathcal{F}_{\bar{K}_{n},\,n^{C}},\|\cdot\|_{\infty}\bigr{)}\leq\log|\mathscr{E}|+\max_{\mathcal{E}}\log N_{B}\lesssim n\varepsilon_{n}^{2}+n\varepsilon_{n}^{2}\lesssim n\varepsilon_{n}^{2}$
This completes the proof.
@@
$B_{n}=\left\{(f,\sigma):\sum_{i=1}^{n}K(p_{0,i},p_{f,\sigma,i})\leq n\epsilon_{n}^{2},\ \sum_{i=1}^{n}V(p_{0,i},p_{f,\sigma,i})\leq n\epsilon_{n}^{2}\right\}$